OpenStudy (anonymous):

A seagull can dive 45 meters below the surface of the water to reach a fish. It dives 11% of the total depth every 1 second. What is the depth of the seagull after 2 seconds?

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

will fan and give medal

OpenStudy (dancergirl45):

what lesson are you in

OpenStudy (anonymous):

what do u mean

OpenStudy (dancergirl45):

nevermind

OpenStudy (dancergirl45):

ok so i think it is going to be 8 seconds, it that one of your choices?

OpenStudy (dancergirl45):

I mean the depth is 8

OpenStudy (anonymous):

nope

OpenStudy (dancergirl45):

What are your choices

OpenStudy (anonymous):

-2.475, - 4.95, -9.9, - 7.425

OpenStudy (michele_laino):

every 1 second the seagull dive for a distance of: \[\frac{{11}}{{100}} \cdot 45 = ...?\]

OpenStudy (dancergirl45):

-4.95

OpenStudy (anonymous):

yip i just got that thanks to all two of u

OpenStudy (dancergirl45):

*2=

OpenStudy (dancergirl45):

-9.9

OpenStudy (michele_laino):

so, after 2 seconds the seagull dives for a distance of \(-4.95 \cdot 2=...\)

OpenStudy (anonymous):

yip

OpenStudy (anonymous):

i have a couple more questions i need help on i will post them